A gas molecule having a speed of 300m/s horizontally collides elastically with another molecule of the same mass is initially at rest. After collision the first molecule moves at an angle of 30° to it's initial direction. Find the speed of each molecule after collision and the direction of the second molecule.
1
Expert's answer
2019-12-24T14:30:44-0500

As per the given question,

Mass of the first molecule=m

Speed of the molecule =300 m/sec

Mass of the second molecule =m

Initial velocity of the second molecule =0 m/sec

After the collision, first molecule is moving at an angle = 30

We have to determine the velocity of each molecule and direction of the second molecule?

Now,

Initial momentum of the molecule "= mv= 300m kg-m\/sec"


"u_1 = v_1\cos30 + v_2\cos\theta_2-------(1)"


"v_1\sin30 = v_2\sin\theta_2----------(2)"


and "u^2_1=v^2_1+v^2_2-----------(3)"


Now, squaring and adding equation (i) and (ii),


"\Rightarrow u^2_1+v^2_1−2u_1v_1\cos30=v^2_2-----------(4)"


Now, from the equation (3) and (4)


"2v^2_1=2u_1v_1\cos30"


from the above,


"v_1=u_1\cos30=260m/s"


"v_2 = 150 m\/s"

from the equation(ii)

"\theta=60^\circ"